# Unraveling hydrodynamization using ultracold 1D gases

### v2

@@ Abstract
We study the quantum evolution of 1D Bose gases immediately after several variants of high-energy quenches, both experimentally and theoretically. Using the advantages conveyed by the relative simplicity of these nearly integrable many-body systems, we are able to differentiate the behavior of two distinct but often temporally overlapping processes, hydrodynamization and local prethermalization. There is a universal character to our findings, which can be applied to the short-time behavior of any interacting many-body quantum system after a sudden high-energy quench. We specifically discuss its potential relevance to heavy-ion collisions.
@@ I Introduction
It is axiomatic in quantum mechanics that the unitary dynamics of a quantum state proceeds due to the relative evolution of the Bohr factors of the unchanging occupied eigenstates [1]. However, it is usually hard to get insight into the way any particular system evolves based on this high level view. Nearly integrable many-body systems are special cases in this regard, since they can be fully characterized in terms of very long lived quasiparticles that arise from interactions among particles. The momenta of these quasiparticles are known as rapidities, the distribution of which fully characterizes equilibrium states [2]. Rapidity distributions can be calculated in experimentally relevant models and they can be measured in 1D gases [3, 4]. In nearly integrable systems, consideration of the Bohr factors associated with quasiparticle states provides deep and surprisingly generic insights into the dynamics that follow a quantum quench.
The quasiparticle description provides the foundation for understanding “hydrodynamization” in nearly integrable gases of ultracold atoms in one dimension (1D) [5]. The term hydrodynamization was first introduced to describe the notably short period of time after the start of a relativistic heavy ion collision before hydrodynamics can be used to describe the time evolution of the system [6, 7, 8, 9, 10, 11, 12]. A characteristic feature of hydrodynamization is that energy gets redistributed among momentum modes that span the many-body system. Local thermalization, in contrast, involves redistribution among nearby momentum modes. The extent to which the two processes are temporally distinct depends on the details of the system and the quench, but it is generally the case that hydrodynamization runs its course before local thermalization is complete.
In Ref. [5], we showed that the time evolution after a large energy quench (implemented via a Bragg pulse) exhibits hydrodynamization before local prethermalization is completed. The quasiparticle picture allowed us to identify two time scales, hydrodynamization and momentum-dependent local prethermalization, which scale with the Bragg momentum differently from each other. We conjectured in Ref. [5] that, since hydrodynamization occurs on very short time scales, far-from-integrable quantum systems that exhibit hydrodynamization may admit a quasiparticle description such as the one we used in near-integrable quantum systems. The quasiparticles, which account for strong correlations and explain the rapid energy transfer among widely separated momentum modes, do not have time to decay during hydrodynamization even if the system is far-from-integrable. We therefore expect that the understanding gained from consideration of nearly integrable 1D gases can be applied to generic interacting quantum systems in the short time immediately following a quench.
The first goal in this work is to explore the conditions necessary for hydrodynamization to occur in “high-energy” quenches, and to differentiate two characteristic time scales related to hydrodynamization: a single hydrodynamization coherence time $T^{c}_{\rm hd}$ identified in Ref. [5], and a hydrodynamization damping time $T^{d}_{\rm hd}$ that depends on the observable being considered. The second goal is to identify the times $T_{\rm lp}$ associated with local prethermalization for different observables, and to explore local prethermalization in the presence and absence of hydrodynamization. We demonstrate, experimentally and theoretically, that the existence of distinct hydrodynamization epochs requires that there be a multimodal quasiparticle energy distribution after the quench. We consider two different quenches, a single Bragg pulse (as in Ref. [5]) and a quantum Newton’s cradle (QNC) sequence [13, 14]. Both quenches generate “high-energy” states with multimodal rapidity distributions, but only the single Bragg pulse generates a multimodal quasiparticle energy distribution. We show, experimentally and theoretically, that the single Bragg pulse leads to hydrodynamization, while the QNC sequence does not.
We then explore theoretically the extent to which a translationally invariant model reproduces the results of Bragg quenches in the presence of the trap. It turns out that the trap is not important, so we explore hydrodynamization and local prethermalization in a translationally invariant system. For that case, we directly write the initial out-of-equilibrium many-quasiparticle state as a finite superposition of plane waves, rather than creating it with Bragg pulses. Our theoretical analysis shows that local prethermalization can occur very rapidly, regardless of whether or not there is hydrodynamization. We put to use the insights gained in this work to provide estimates for the hydrodynamization times and the fastest local prethermalization times in heavy-ion collisions.
The presentation is organized as follows. In Sec. II, we introduce the experimental setup and review its theoretical modeling. We also review our overall understanding of the quantum dynamics following a single Bragg pulse in a homogeneous Tonks-Girardeau gas. In Sec. III, we study the time evolution after a single Bragg pulse and after a QNC sequence, in both the experiment and the Tonks-Girardeau gas theory. We then explore theoretically, in Sec. IV, the effect on the hydrodynamization and local prethermalization time scales of density inhomogeneity due to the trapping potential and of finite Bragg pulse duration. In Sec. V, we use finite superpositions of plane waves in a translationally invariant model to explore in detail the necessary conditions for hydrodynamization and local prethermalization. We also determine how the local prethermalization rate for the momentum distribution depends on the specific superpositions considered. The potential relevance of our findings to heavy-ion collisions is discussed in Sec. VI. We summarize our results, and highlight open questions in Sec. VII.

### v3

# Timescales and necessary conditions for hydrodynamization in one-dimensional Bose gases

@@ Abstract
We study the quantum evolution of one-dimensional Bose gases immediately after several variants of high-energy quenches, both theoretically and experimentally. Using the advantages conveyed by the relative simplicity of these nearly integrable many-body systems, we are able to differentiate the behaviors of two distinct but often temporally overlapping processes, hydrodynamization and local prethermalization. We show that the hydrodynamization epoch is itself characterized by two independent timescales, an oscillation period and an observable-dependent damping time. We also show how the existence of a hydrodynamization epoch depends on the exact nature of the high-energy quench. There is a universal character to our findings, which can be applied to the short-time behavior of any interacting many-body quantum system after a sudden high-energy quench. We specifically discuss its potential relevance to heavy-ion collisions.
@@ I Introduction
The term “hydrodynamization” was introduced to describe the notably short period of time between the start of a relativistic heavy-ion collision and when hydrodynamics describes its time evolution [1, 2, 3, 4, 5, 6, 7]. The observed hydrodynamization time is shorter than the expected local thermalization time, which is the time required for the system to reach thermal equilibrium within a hydrodynamic cell. Based on experimental results with nearly integrable gases of ultracold atoms in one dimension and the associated theory, we argued in Ref. [8] that immediately after a high-energy quench the time evolution of any many-body quantum system can be qualitatively understood in a universal way. There is an initial epoch, which by analogy to the relativistic heavy-ion usage we called “hydrodynamization,” during which energy is redistributed across distant momentum modes and short-distance density fluctuations damp out. The second, partially temporally overlapping epoch is local prethermalization, which is analogous to thermalization in far-from-integrable systems and is characterized by a redistribution of energy among nearby momentum modes. Hydrodynamization and local prethermalization are quantum mechanical phenomena, yet we showed in the experiments in Ref. [8] that generalized hydrodynamics (a classical description) [9, 10, 11, 12, 13] can be used to describe the system before local prethermalization is complete as long as the hydrodynamization epoch is over.
In this paper, again using nearly integrable one-dimensional (1D) gases as a model, we significantly broaden our understanding of hydrodynamization. In addition to the universal hydrodynamization coherence time that we identified in Ref. [8], we identify a new observable-dependent hydrodynamization damping time that sets the lower limit for the application of hydrodynamic descriptions. We show, in both theory and experiments, that the existence of a distinct hydrodynamization epoch requires that there be a multimodal energy distribution after the quench. We further show that, although the extent of temporal overlap of hydrodynamization damping (when it exists) and local prethermalization depends on the system and the observable being considered, the two timescales scale with initial state parameters in the same way. We are able to apply these findings to a quantitative consideration of relativistic heavy-ion collisions. Our estimates of the shortest time after which hydrodynamics can describe the dynamics in those systems are in line with the observations of those collisions.
To see how a nearly integrable system can give insights into nonintegrable system dynamics, it is useful to start with a very high level view of dynamics in quantum systems. The time-evolving wave function $\Psi(\textbf{q},t)$ describing an isolated quantum system can be written as
$\Psi(\textbf{q},t)=\sum_{j}\,c_{j}\,\psi_{j}(\textbf{q})\,e^{-iE_{j}t/\hbar},$ (1)
where q includes the positions of all the particles, $\psi_{j}(\textbf{q})$ are the energy eigenfunctions with eigenenergies $E_{j}$, and the (time independent) $c_{j}$ are the overlaps of the initial state with $\psi_{j}(\textbf{q})$. Unitary dynamics is the result of the relative evolution of the phases $E_{j}t/\hbar$ between the energy eigenfunctions. Equation (1) contains all the system’s dynamical information but it is not useful for large, generic interacting quantum systems, where none of its elements can be calculated. Without explicit calculations, the equation generally provides no way to extract the timescales of processes such as hydrodynamization and local prethermalization.
@@
Nearly integrable many-body systems, however, are special in this regard. They can be characterized in terms of very long lived quasiparticles that arise from interactions among particles. The momenta of these quasiparticles are known as rapidities $\theta$, the distribution of which fully characterizes equilibrium states in the integrable limit [14]. Rapidity distributions, $f(\theta)$, can be calculated in experimentally relevant models and can be measured in 1D gases [15, 16, 17, 18]. Consideration of the relative evolution of the phases associated with quasiparticle states allows one to identify important out-of-equilibrium features, including the timescales associated with hydrodynamization and local prethermalization. Notably, such features can be identified even when exact calculations are out of reach. We conjectured in Ref. [8] that, shortly after high-energy quenches, far-from-integrable quantum systems admit a similar description that incorporates interactions into quasiparticles, with the distinction that they are short-lived and likely hard to calculate. But since the hydrodynamization epoch is so short, quasiparticles are not expected to decay in that time. We therefore expect that the expanded understanding of hydrodynamization we develop here from consideration of nearly integrable 1D gases can be applied to generic interacting quantum systems, as we do here for relativistic heavy-ion collisions.
The presentation is organized as follows. In Sec. II, we introduce the experimental setup and review its theoretical modeling. We also present our overall understanding of the quantum dynamics that follow a single Bragg pulse acting on a Tonks-Girardeau gas. In Sec. III, we theoretically study hydrodynamization and local prethermalization in a translationally invariant system with a trimodal rapidity distribution. Changing the width of the rapidity peaks and their separation allows us to probe, for the first time, the two timescales associated with hydrodynamization as well as the local prethermalization time for the momentum distribution. In Sec. IV, we consider the case of a symmetric bimodal rapidity distribution, which allows us to demonstrate the absence of hydrodynamization in high energy quenches that lack multimodal energy distributions. The latter conclusion is tested experimentally in Sec. V using a quantum Newton’s cradle [19] initial state. The relevance of our findings to heavy-ion collisions is discussed in Sec. VI. We summarize our results, and highlight open questions in Sec. VII.